OpenStudy (anonymous):

How can x^2 + 3x + 1 = 2x^2 + 2x + 3 be set up as a system of equations?

OpenStudy (loser66):

don't understand the question

OpenStudy (anonymous):

how can I make that two separate equations

OpenStudy (anonymous):

subtract one from the other and end up with a quadratic equation

OpenStudy (anonymous):

\[x^2 + 3x + 1 = 2x^2 + 2x + 3\\ \iff x^2-x+2=0 \]

OpenStudy (anonymous):

y = x^2 + 3x + 1 y = 2x^2 + 2x + 3 y = 2x^2 y = x^2 + 2x + 3 y = 2x^2 + 2x y = 2x^2 + 9 y = x^2 y = 2x^2 + 2x + 3

OpenStudy (loser66):

@topstryker I am with you

OpenStudy (loser66):

chose the first one

OpenStudy (anonymous):

are you sure??

OpenStudy (loser66):

@satellite73 gives us the simplified form of the problem while the question is "separate them" I think you are correct.

OpenStudy (loser66):

what grade are you?

OpenStudy (anonymous):

10

OpenStudy (loser66):

which course/

OpenStudy (anonymous):

algebra 2